OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 2x - one over four of (12x + 4) = 3 Step 1: 2x + 3x + 1 = 3 Step 2: 5x + 1 = 3 Step 3: 5x + 1 - 1 = 3 - 1 Step 4: 5x = 2 In which step was the error made and why? Step 1, because incorrect signs have been used Step 2, because the terms have been added incorrectly Step 3, because a number has been subtracted from both sides of the equation No error was made

OpenStudy (anonymous):

look at the original equation and write it out, it should be obvious once you write in numbers

OpenStudy (anonymous):

Yea, Make sure you read your math lessons because they show you how to do this.

OpenStudy (anonymous):

Well...what did you get?

OpenStudy (anonymous):

im still working on it im doing two classes at once

OpenStudy (anonymous):

okay just tell me what you get...

OpenStudy (anonymous):

Thankyou for being patient

OpenStudy (anonymous):

It's fine I'm here to help.

OpenStudy (anonymous):

I think c

OpenStudy (anonymous):

hopefully

OpenStudy (anonymous):

Okay...SHow me were you solved that...

OpenStudy (anonymous):

like~show your steps...

OpenStudy (anonymous):

I looked at it and your supposed to replace the subtraction with addition i think so i said c

OpenStudy (anonymous):

Hold on...

OpenStudy (anonymous):

i just looked at step 3

OpenStudy (anonymous):

I don't think it's step 3.

OpenStudy (anonymous):

\[2x - \frac{ 1 }{ 4 }(12x + 4) = 3\] here is your equation

OpenStudy (anonymous):

Look it over again...

OpenStudy (anonymous):

sTEP 1 BECAUSE HE DIDNT DO PERINTHESIS FIRST

OpenStudy (anonymous):

Ummm, I guess go for it...I can't tell you the answer! Besides the moderator that got me the first time is viewing this.

OpenStudy (anonymous):

ok thankyou for yalls time i appreciate it.

OpenStudy (anonymous):

No problem...Bye Compassionate!

OpenStudy (debbieg):

You started with this: \[2x - \frac{ 1 }{ 4 }(12x + 4) = 3\] In step 1, he got this: \[2x +3x + 1 = 3\] The error is in step 1, but it isn't that he "didn't do the parentheses".

OpenStudy (anonymous):

oh i got it right then thanks yall.